Venom owns land with an adjusted basis of S280,000 and a fair market value of S150,000. He sells the land for $115,000 to Phenom, Inc., a corporation in which he owns 60% of the stock. Determine the amount of realized and recognized gain or loss to Venom and the adjusted basis for Phenom. (5 points) 3. 4. In 2018, Riot purchased 200 shares of Misanthropy Corporation stock for $24,000 on January 1, 2015. He sells 50 shares of the 200 shares purchased on January 1, 2015, for S5,400. Twenty-five days earlier, he had purchased 60 shares for S5,800. What is Riots realized and recognized gain or loss on the sale of the stock, and what is his basis in the 60 shares purchased 25 days earlier? (10 points)
Answer #1

3. the loss from the transaction is $35,000 (the difference between $150,000 - 115,000)

and the ownership is 60%, hence his loss will be= 35,000* 0.60= $21,000

4. the average buy price is= 24,000/200= $120

average sell price of 50 stocks= 5800/50= $116

as he sold of 50 shares, his loss will be= $4*50= $200